OpenStudy (anonymous):

The root x = 1 has multiplicity 2 for the function f(x)= x^3 - x^2 - x + 1 A. True B. False I think it is True

OpenStudy (anonymous):

How many times can you factor out x-1 from the expression?

OpenStudy (anonymous):

f(x)= x^3 - x^2 - x + 1 = x^2(x - 1) - (x - 1) = (x^2 - 1)(x - 1) = (x + 1)(x - 1)(x - 1)